IP angles scan experiment - 05 March 2007

LER Single beam: Push current

TimeComments
17:17 Trickling up the LER to 600 mA
17:23 I_LER = 600 mA. Uli changed the gain of the LFB. SCP died: saved file: LERIS_1_7MAR
17:25 I_LER=700 mA. SCP died again: saved file: LERIS_2_7MAR. LINAC down.
17:55 LINAC back up. I_LER=600 mA
18:10 I_LER = 900 mA. Angle check: X-prime=9.1 mrad, Y-prime=-1.37
18:20 I_LER=1090 mA
18:30 I_LER=1300 mA
18:40 I_LER=1480 mA. ST BPMs are saturating. Stopping here. File: LERIS_3_7MAR